Two planes leave Pittsburgh and Philadelphia at the same time, each going to the other city. One plane flies 25 miles per hour faster than the other. Find the air speed of each plane if the cities are 275 miles apart and the planes pass each other after 40 minutes of flying time.
Answer by mananth(16949)

Two planes leave Pittsburgh and Philadelphia at the same time, each going to the other city. One plane flies 25 miles per hour faster than the other. Find the air speed of each plane if the cities are 275 miles apart and the planes pass each other after 40 minutes of flying time.
...
the planes are flying towards each other
one plane speed = xmph
other plane speed = x+25
..
effective speed = x+25+x= 2x+25 mph
...
40 minutes = 2/3 hours
time = distance / speed
275/(2x+25)=2/3
..
3*275 = 2(2x+25)
825=4x+50
825-50=4x
775=4x
x= 775/4
193.75 mph speed of one plane.
193.75+25 =218.75mph = speed of other plane
...
m.ananth@hotmail.ca